Syllabus for MATH 557, Fall 2013

Stephen G. Simpson Pennsylvania State University www.math.psu.edu/simpson/courses/math557

August 23, 2013

We will cover Chapters 1–5 of the Mathematical Logic notes, Chapters 1 and 2 of the Foundations of Mathematics notes, and all of the Incompleteness and Undecidability notes.

August 27 - ML notes: §§1.1–1.3 August 29 - §1.4 September 3 - §1.5 September 5 - §§1.6–1.8 September 10 - §§2.1–2.2 September 12 - §2.3 September 17 - §2.4 September 19 - §2.5 September 24 - §2.6 September 26 - §2.7 October 1 - §3.1, §3.3 except completeness of LH October 3 - §3.2 plus completeness of LH October 8 - §3.4 October 10 - §3.5 October 15 - §§4.1–4.2 October 17 - §§4.3–4.5 October 22 - Midterm Exam October 24 - ML notes: §§5.1–5.2 October 29 - §5.3 October 31 - §5.4, §5.8 November 5 - §5.5 November 7 - §§5.6–5.7 November 12 - FoM notes: §§1.1–1.2 November 14 - §1.3 November 19 - §§1.4–1.5 November 21 - §§1.6–1.8 December 3 - §§2.1–2.3 December 5 - I&U notes: §1, first half December 10 - §1, second half December 12 - §2